### v1

@@ 5.4.1. Qualitative Data
For qualitative data that includes in-study reflections and interview transcripts, we conducted reflexive thematic analysis (Braun and Clarke, 2019). The analysis was led by a graduate student formally trained in qualitative methods, with prior research experience across HCI domains, including human-centered AI, computing education, and visualization. The process was collaborative, involving iterative discussions on codes and themes with a broader research team that included two faculty supervisors and other HCI and visualization researchers. As part of our reflexive practice, we continually questioned how our backgrounds might shape our interpretations. For instance, as designers and researchers of interactive systems, we asked: How do our own writing and feedback-seeking practices influence what we notice in participants’ accounts? While these reflexive discussions do not eliminate subjectivity, they were essential for ensuring our workflow remained an interpretative and dialogic process.
@@ 5.4.2. Quantitative Data
To supplement our primary qualitative data from interviews and observations, we administered short surveys both before and immediately after the study. Pre-study survey includes questions on demographics and individuals’ self-reported reflection practices using their own habitual QDA tools. The post-study survey included Likert-scale items assessing the experience and usability of the tool, as well as questions about the frequency of reflection using Reflexis. For the frequency question, we paired the responses from each participant from the pre-study survey and post-study survey to examine any differences in analysis practices.
@@ 5.4.3. LLM Technical Analysis
While the ultimate utility of an interpretation in RTA is determined by the researcher, we sought to verify the general quality of the output from LLM-enabled features through a two-stage process: a preliminary quality check during the pilot study and a post-hoc technical analysis. During the pilot phase, we iteratively tested and refined the system prompts to ensure baseline coherence. Following the evaluation study, two co-authors performed a post-hoc analysis of the interaction logs, reviewing every occurrence of Code Drift Alert ($N=7$) and Positionality-aware Discussion Prompt ($N=18$) as they naturally revealed during the study.
@@ 5.5. Rationale and Limitations
A single-condition design privileges depth of interpretive work over tool-to-tool comparisons. This limits claims about comparative effectiveness but aligns with RTA’s interpretivist commitments and avoids incorporating non-standard “baselines”  (Braun and Clarke, 2023; Carminati, 2018), as discussed in Section 5.2.1. We partially mitigate the lack of a baseline by (a) eliciting participants’ customary tool use and practices, and (b) observing disagreement-focused collaboration during the study session. Using an LLM-generated dataset from a real-world example preserves the original civic context while minimizing sensitivity risks (Tsai et al., 2016); however, synthetic data may narrow variability relative to field data and could subtly steer theme salience.
6. User Study Result
6.1. Reframing Reflexivity: Reflexis Encourages Granular, In-Situ Reflection Amidst Workflow Tensions (RQ1)
Figure 5. Reflexis: Change in self-reported frequency of in-situ reflexivity. The Sankey diagram visualizes how participants perceived their practice of pausing to write a rationale for a code choice, comparing their recalled frequency from their typical workflows (“Before”) with their reported frequency during the study using Reflexis (“After”). This data is presented to illustrate a perceived change in practice, not as a direct measure of behavior.
Integrating structured, lightweight reflection directly into the coding interface via the ReflexiveLens prompt a perceived shift in participants’ analytical practices, from high-level, delayed habits toward more granular, in-situ engagement with their data. To illuminate this perceived shift, we asked participants to self-report the frequency of their in-situ reflection for both their typical workflows and their work during the study. While we acknowledge that comparing recalled behavior to in-the-moment experience is not a controlled measure, it provides insight into the participants’ own sense of their changing practice. As illustrated in Figure 5, this self-reported data shows a notable trend toward more frequent reflection when using Reflexis.
@@ 6.1.1. Providing a New, Integrated Space for Situated Reflection
A primary benefit of ReflexiveLens was its ability to address the critical challenge of disconnected and unorganized reflective workflows. Participants frequently compared the tool’s integrated nature favorably against their current manual methods. $P_{11}$, for example, described her typical process of keeping a “side document” for reflections, which she found “really difficult to, be going back and forth, like, trying to link excerpts” and made her prone to losing context over time.
@@ 6.1.2. Shifting Reflection from a Broad, Delayed Act to a Deliberate Practice
The most novel aspect of ReflexiveLens was its ability to facilitate reflection at the snippet level, a significant departure from most participants’ existing habits of reflecting broadly on an entire interview “at the end of a coding session” ($P_{7}$). For instance, we observed $P_{7}$ use the ReflexiveLens prompt immediately after applying an ‘emotional response‘ code to a passage describing an ‘angry’ and ‘apologetic’ resident. In her reflection, she articulated a direct link between her coding choice and her positionality as a woman. This behavior presents how the tool transformed a typically delayed, high-level task into an immediate, code-specific, and deeply personal analytical act.
The tool’s ability to support reflection on specific codes as they were applied was considered a valuable improvement. This granular, in-the-moment approach was not just a siloed activity; it also scaffolded collaborative sensemaking. $P_{12}$ recalled how seeing their teammate’s ($P_{11}$) reflexive notes prompted them to “thinking very hard” about their reasoning, thereby encouraging deeper engagement with both the data and their own interpretive process ($P_{12}$).
@@ 6.1.3. Preferences in Reflective Scope: A Desire for Higher-Level, Networked Memos
While the availability to reflect on a granular level was highly valued, the tool’s per-passage focus did not perfectly align with the workflow of all participants. Some participants ($P_{1}$, $P_{5}$, $P_{6}$, $P_{9}$), described their most critical reflections as occurring at a higher level of abstraction. $P_{6}$ found it difficult to be meaningfully reflexive on isolated passages, noting, “I think of it more in terms of how I apply multiple codes over time.” $P_{9}$, whose memoing process typically happens at the theme-level, expressed a need for features that support relational analysis:
@@ 6.2.1. The Analysis History Was Universally Embraced as a Mechanism for Transparency and Rigor
The Analysis History feature was met with strong enthusiasm across all participant pairs. This positive reception was confirmed in our survey, where 11 of 12 participants agreed that the feature increased the transparency of their research process and allowed them to easily see code changes over time. The feature’s primary value, as described by participants, was its ability to automate and centralize a process that is currently manual, fragmented, and tedious. Researchers described their ad-hoc workflows involving “a Google slideshow, my Atlas.ti, and a couple Google Doc memos” ($P_{5}$), or tracking changes via the “clunky” edit history in Google Sheets ($P_{6}$). The system’s ability to provide a visual and integrated history was seen as a significant improvement. $P_{12}$, whose team currently tracked changes manually in an Excel sheet, saw it as a powerful automation of a necessary but laborious task.
Participants identified two high-impact use cases for this newfound transparency. First, researchers connected the feature directly to the task of writing up their methodology. $P_{3}$ explained that the visual graph is a “very useful way to think back and see how different codes merged or split up,” which would help in reporting the analytical process. $P_{4}$ described the feature as an audit trail for their work:
“I think it’s definitely going to be helpful, for transparency’s sake. It’s like you hav a blockchain or something equivalent, so you have like, a universally agreed upon, history of changes.” ($P_{4}$)
@@ 6.2.2. The Code Drift Alert Sparked Critical Reflection, Highlighting the Value of Proactive AI Partnership
Despite the mixed feedback on its timing, participants demonstrated sustained engagement after the alert was shown. This was visible even when a participant was taken aback. For instance, after an alert “took [her] by surprise,” $P_{2}$ was observed immediately pausing to refine her code’s definition. $P_{1}$’s actions further illustrated this pattern of deliberate engagement. She described accepting the alert’s suggestion to create a new code on one occasion, while on another, she confidently overrode it.
@@ 6.3. Scaffolding Collaborative Interpretation by Focusing and Framing Disagreement (RQ3)
To understand how our system scaffolds collaborative interpretation (RQ3), we analyzed how research pairs engaged with features designed to surface and contextualize interpretive differences. The findings show that the Discussion Focus was seen as a practical enhancement to existing workflows. The Positionality-aware Discussion Prompt, although a more context-dependent feature, was recognized for its unique potential to foster deeper, methodologically-grounded reflexive dialogue.
@@
Participants immediately identified the Discussion Focus features as a practical and powerful tool for structuring collaboration. This was one of the most highly-rated features in our survey, with 11 of 12 participants agreeing it helped them notice meaningful disagreements and made it significantly easier to start or engage in discussion with collaborators. This was evident in pairs’ immediate adoption of the feature to structure their conversation. For instance, at the start of their discussion, $P_{2}$ announced, “I just toggled on [the show diff view toggle]… so we’re just gonna focus on that,” while $P_{11}$ began by acting on the discussion status icon’s information: “I’m gonna start with the one that says, needs discussion.”
@@ 6.4. Overall User Experience
Beyond the findings for specific features, our study revealed a strong positive reception of Reflexis as a complete system, positioning it as a significant improvement over participants’ existing qualitative analysis workflows. Which was supported by the unanimous agreement from participants on the system’s core principles: every single participant (12/12) reported feeling “in control of analytic decisions” and agreed the system “supported an iterative research process.”
Participants frequently attributed this holistic improvement to the tool’s integrated nature, which removes the cognitive friction of their current, fragmented workflows. $P_{7}$, for instance, contrasted Reflexis favorably with their “document-heavy” Google Docs method. The tool’s ability to facilitate collaboration was a standout benefit, addressing a critical gap in the current software landscape. As $P_{5}$ explained:
@@
Of course, the experience was not without limitations. Some participants noted the study’s time constraints created a feeling of pressure, preventing them from the deeper, more time-intensive engagement the tool affords ($P_{9}$). The most common requests for improvement were visual, including color-coded highlights and a hierarchical codebook to better organize themes ($P_{7},P_{11},P_{12}$).
We report the detailed descriptive statistics in Appendix A.5.
@@ 7. Discussion: Designing for Deliberation in Human-AI Systems to Support Human Collaboration
As AI models increasingly automate complex cognitive tasks, the challenge for HCI is shifting from designing efficient human-AI pipelines to a more nuanced goal: designing for deliberation. This approach, which builds on calls to treat ambiguity as a design resource (Gaver et al., 2003) and to create technologies that support reflection (Baumer, 2015), prioritizes systems that intentionally support the deep, and often contentious, processes of human interpretation (Ma et al., 2025). It stands in contrast to an uncritical focus on automation, which risks diminishing creative originality and suppressing critical thought (Kumar et al., 2025; Lee et al., 2025).
@@ 7.1. Augmenting Agency: Shifting from Automation to Deliberation
This approach addresses emerging concerns regarding the epistemic risks of over-automating cognitive work. Recent research indicates that reliance on generative models can lead to the “homogenization” of creative outputs (Kumar et al., 2025), while AI-generated explanations may paradoxically suppress critical thinking and promote overreliance (Lee et al., 2025; Kim et al., 2025). In qualitative analysis, where the scientific value is derived from the depth of the researcher’s interpretive effort, such automation is particularly counterproductive (Dourish, 2014; Jowsey et al., 2025). Reflexis mitigates these risks by casting the AI as a non-prescriptive observer. For instance, the Code Drift Alert does not “correct” the researcher but rather prompts potential concept drift to force a deliberate choice. Participants reported that this friction made them “more precise” and “self-aware”($P_{11}$), illustrating how AI can be designed to deepen, rather than bypass, intellectual rigor.
This design pattern, which employs AI to encourage deeper immersion of context, surface latent tensions, encourage reasoned deliberation, and even demand human-led critical inquiry, has broad implications for other domains of interpretive and creative work: For Scholarly and Creative Writing $\Rightarrow$ An AI system could act as a developmental editor by tracking the evolution of arguments or character arcs in a long-form document. It would highlight potential inconsistencies or tonal shifts for the author’s review, preserving their narrative authority while augmenting their capacity for self-critique, designed to deepen, rather than bypass, intellectual rigor. For Design and Engineering $\Rightarrow$ An AI could juxtapose a designer’s specifications against their stated goals or established design patterns. It could then surface tensions, for example, a conflict between a requirement for “ease of use” and an architectural choice that adds complexity. This prompts the designer to explicitly rationalize trade-offs early in the implementation process.
@@ 7.2. Designing for Analytical Provenance to Support Resumption and Longer-term Usage
Reflexis is designed to bridge this gap by prioritizing a meaningful, auditable narrative of conceptual development over an exhaustive interaction log. We operationalize this via data-level provenance – which visualizes the structural evolution of analytic actions (e.g., splits, merges, renames) (Viégas et al., 2004), augmented by lightweight rationale provenance – the brief reflections and justifications linked to individual snippets. Our findings (RQ2) suggest that this combination effectively balances utility and manageability. By transforming the codebook from a static list into a living history, the system shows potential as an external memory scaffold and is “a very useful way to think back.” This allows researchers to combat the “drift” and fatigue inherent in longitudinal work (O’Connor and Joffe, 2020). While longitudinal deployment is necessary to confirm long-term adoption, this immediate utility suggests that provenance views within Reflexis can facilitate re-immersion and transparent reporting, without the overhead of manual journaling.

### v2

@@ 5.4.1. Qualitative Data
We conducted RTA for qualitative data including in-study reflections and interview transcripts (Braun and Clarke, 2019). The analysis was led by a graduate student formally trained in qualitative methods, with prior research experience across HCI domains, including human-centered AI, computing education, and visualization. The process was collaborative, involving iterative discussions on codes and themes with a broader research team that included two faculty supervisors and other HCI and visualization researchers. As part of our reflexive practice, we continually questioned how our backgrounds might shape our interpretations. For instance, as designers and researchers of interactive systems, we asked: How do our own writing and feedback-seeking practices influence what we notice in participants’ accounts? While these reflexive discussions do not eliminate subjectivity, they were essential for ensuring our workflow remained an interpretive and dialogic process.
@@ 5.4.2. Quantitative Data
To supplement our primary qualitative data from interviews and observations, we administered short surveys both before and immediately after the study. The pre-study survey included questions on demographics and individuals’ self-reported reflection practices using their own habitual QDA tools. The post-study survey included Likert-scale items assessing the experience and usability of the tool, as well as questions about the frequency of reflection using Reflexis. For the frequency question, we paired the responses from each participant from the pre-study survey and post-study survey to examine any differences in analysis practices.
@@ 5.4.3. LLM Technical Analysis
While the ultimate utility of an interpretation in RTA is determined by the researcher, we sought to verify the general quality of the output from LLM-enabled features through a two-stage process: a preliminary quality check during the pilot study and a post-hoc technical analysis. During the pilot phase, we iteratively tested and refined the system prompts to ensure baseline coherence. Following the evaluation study, two co-authors performed a post-hoc analysis of the interaction logs, reviewing every occurrence of Code Drift Alert ($N=7$) and Positionality-aware Discussion Prompt ($N=18$) as they naturally triggered during the study.
@@ 5.5. Rationale and Limitations
A single-condition design privileged depth of interpretive work over tool-to-tool comparisons. This limits claims about comparative effectiveness but aligns with RTA’s interpretivist commitments and avoids incorporating non-standard “baselines”  (Braun and Clarke, 2023; Carminati, 2018), as discussed in Section 5.2.1. We partially mitigated the lack of a baseline by (a) eliciting participants’ customary tool use and practices, and (b) observing disagreement-focused collaboration during the study session. Using an LLM-generated dataset from a real-world example preserves the original civic context while minimizing sensitivity risks (Tsai et al., 2016); however, synthetic data may narrow variability relative to field data and could subtly steer theme salience.
6. User Study Results
6.1. Reframing Reflexivity: Reflexis Encouraged Granular, In-Situ Reflection Amidst Workflow Tensions (RQ1)
Figure 5. Reflexis: Change in self-reported frequency of in-situ reflexivity. The Sankey diagram visualizes how participants perceived their practice of pausing to write a rationale for a code choice, comparing their recalled frequency from their typical workflows (“Before”) with their reported frequency during the study using Reflexis (“After”).
Integrating structured, lightweight reflection directly into the coding interface via the ReflexiveLens prompted a perceived shift in participants’ analytical practices, from high-level, delayed habits toward more granular, in-situ engagement with their data. To investigate this perceived shift, we asked participants to self-report the frequency of their in-situ reflection for both their typical workflows and their work during the study. While we acknowledge that comparing recalled behavior to in-the-moment experience is not a controlled measure, it provides insight into the participants’ own sense of their changing practice. As illustrated in Figure 5, this self-reported data shows a notable trend toward more frequent reflection when using Reflexis.
@@ 6.1.1. Providing a New, Integrated Space for Situated Reflection
A primary benefit of ReflexiveLens was its ability to address the critical challenge of disconnected and unorganized reflective workflows. Participants frequently compared the tool’s integrated nature favorably against their current manual methods. $P_{11}$, for example, described her typical process of keeping a “side document” for reflections, which she found “really difficult to be going back and forth, like, trying to link excerpts” and made her prone to losing context over time.
@@ 6.1.2. Shifting Reflection from a Broad, Delayed Act to a Deliberate Practice
The most novel aspect of ReflexiveLens was its ability to facilitate reflection at the snippet level, a significant departure from most participants’ existing habits of reflecting broadly on an entire interview “at the end of a coding session” ($P_{7}$). For instance, we observed $P_{7}$ use the ReflexiveLens prompt immediately after applying an ‘emotional response’ code to a passage describing an “angry” and “apologetic” resident. In her reflection, she articulated a direct link between her coding choice and her positionality as a woman. This behavior illustrated how the tool transformed a typically delayed, high-level task into an immediate, code-specific, and deeply personal analytical act.
The tool’s ability to support reflection on specific codes as they were applied was considered a valuable improvement. This granular, in-the-moment approach was not just a siloed activity; it also scaffolded collaborative sensemaking. $P_{12}$ recalled how seeing their teammate’s ($P_{11}$) reflexive notes prompted them to “think very hard” about their reasoning, thereby encouraging deeper engagement with both the data and their own interpretive process ($P_{12}$).
@@ 6.1.3. Preferences in Reflective Scope: A Desire for Higher-Level, Networked Memos
While the ability to reflect on a granular level was highly valued, the tool’s per-passage focus did not perfectly align with the workflow of all participants. Some participants ($P_{1}$, $P_{5}$, $P_{6}$, $P_{9}$) described that their most critical reflections occur at a higher level of abstraction. $P_{6}$ found it difficult to be meaningfully reflexive on isolated passages, noting, “I think of it more in terms of how I apply multiple codes over time.” $P_{9}$, whose memoing process typically happened at the theme-level, expressed a need for features that support relational analysis:
@@ 6.2.1. The Analysis History Was Universally Embraced as a Mechanism for Transparency and Rigor
The Analysis History feature was met with strong enthusiasm across all participant pairs. This positive reception was confirmed in our survey, where 11 of 12 participants agreed that the feature increased the transparency of their research process and allowed them to easily see code changes over time. The feature’s primary value, as described by participants, was its ability to automate and centralize a process that is currently manual, fragmented, and tedious. Researchers described that their ad-hoc workflows involved “a Google slideshow, my Atlas.ti, and a couple Google Doc memos” ($P_{5}$), or tracking changes via the “clunky” edit history in Google Sheets ($P_{6}$). The system’s ability to provide a visual and integrated history was seen as a significant improvement. $P_{12}$, whose team currently tracked changes manually in an Excel sheet, saw it as a powerful automation of a necessary but laborious task.
Participants identified two high-impact use cases for this newfound transparency. First, researchers directly connected the feature to the task of writing up their methodology. $P_{3}$ explained that the visual graph is a “very useful way to think back and see how different codes merged or split up,” which would help in reporting the analytical process. $P_{4}$ described the feature as an audit trail for their work:
“I think it’s definitely going to be helpful, for transparency’s sake. It’s like you have a blockchain or something equivalent, so you have like, a universally agreed upon, history of changes.” ($P_{4}$)
@@ 6.2.2. The Code Drift Alert Sparked Critical Reflection, Highlighting the Value of Proactive AI Partnership
[…]
@@ 6.3. Scaffolding Collaborative Interpretation by Focusing and Framing Disagreement (RQ3)
To understand how our system scaffolds collaborative interpretation (RQ3), we analyzed how research pairs engaged with features designed to surface and contextualize interpretive differences. The findings showed that the Discussion Focus was seen as a practical enhancement to existing workflows. The Positionality-Aware Discussion Prompt, although a more context-dependent feature, was recognized for its unique potential to foster deeper, methodologically-grounded reflexive dialogue.
@@
Participants immediately identified the Discussion Focus features as a practical and powerful tool for structuring collaboration. This was one of the most highly rated features in our survey, with 11 of 12 participants agreeing that it helped them notice meaningful disagreements and made it significantly easier to start or engage in discussions with collaborators. This was evident in pairs’ immediate adoption of the feature to structure their conversation. For instance, at the start of their discussion, $P_{2}$ announced, “I just toggled on [the show diff view toggle]… so we’re just gonna focus on that,” while $P_{11}$ began by acting on the discussion status icon’s information: “I’m gonna start with the one that says, needs discussion.”
@@ 6.4. Overall User Experience
Beyond the findings for specific features, our study revealed a strong positive reception of Reflexis as a complete system, positioning it as a significant improvement over participants’ existing qualitative analysis workflows. This was supported by the unanimous agreement from participants on the system’s core principles: every single participant (12/12) reported feeling “in control of analytic decisions” and agreed the system “supported an iterative research process.”
Participants frequently attributed this holistic improvement to the tool’s integrated nature, which removed the cognitive friction of their current, fragmented workflows. $P_{7}$, for instance, contrasted Reflexis favorably with their “document-heavy” Google Docs method. The tool’s ability to facilitate collaboration was a standout benefit, addressing a critical gap in the current software landscape. As $P_{5}$ explained:
@@
Of course, the experience was not without limitations. Some participants noted the study’s time constraints created a feeling of pressure, preventing them from the deeper, more time-intensive engagement the tool affords ($P_{9}$). The most common requests for improvement were visual, including color-coded highlights and a hierarchical codebook to better organize themes ($P_{7},P_{11},P_{12}$). We report the detailed descriptive statistics in Appendix A.5.
@@ 7. Discussion: Designing for Deliberation in Human-AI Systems to Support Human Collaboration
As AI models increasingly automate complex cognitive tasks, the challenge for HCI is shifting from designing efficient human-AI pipelines to a more intentional goal: designing for deliberation. This approach, which builds on calls to treat ambiguity as a design resource (Gaver et al., 2003) and to create technologies that support reflection (Baumer, 2015), prioritizes systems that intentionally support the deep, and often contentious, processes of human interpretation (Ma et al., 2025). It stands in contrast to an uncritical focus on automation, which risks diminishing creative originality and suppressing critical thought (Kumar et al., 2025; Lee et al., 2025).
@@ 7.1. Augmenting Agency: Shifting from Automation to Deliberation
This approach addresses emerging concerns regarding the epistemic risks of over-automating cognitive work. Recent research indicates that reliance on generative models can lead to the “homogenization” of creative outputs (Kumar et al., 2025), while AI-generated explanations may paradoxically suppress critical thinking and promote overreliance (Lee et al., 2025; Kim et al., 2025). In qualitative analysis, where the scientific value is derived from the depth of the researcher’s interpretive effort, such automation is particularly counterproductive (Dourish, 2014; Jowsey et al., 2025). Reflexis mitigates these risks by casting the AI as a non-prescriptive observer. For instance, the Code Drift Alert does not “correct” the researcher but rather flags potential concept drift to force a deliberate choice. Participants reported that this friction made them “more precise” and “self-aware”($P_{11}$), illustrating how AI can be designed to deepen, rather than bypass, intellectual rigor.
This design pattern, which employs AI to encourage deeper immersion of context, surface latent tensions, allow reasoned deliberation, and even demand human-led critical inquiry, has broad implications for other domains of interpretive and creative work: For Scholarly and Creative Writing $\Rightarrow$ An AI system could act as a developmental editor by tracking the evolution of arguments or character arcs in a long-form document. It would highlight potential inconsistencies or tonal shifts for the author’s review, preserving their narrative authority while augmenting their capacity for self-critique, designed to deepen, rather than bypass, intellectual rigor. For Design and Engineering $\Rightarrow$ An AI could juxtapose a designer’s specifications against their stated goals or established design patterns. It could then surface tensions, for example, a conflict between a requirement for “ease of use” and an architectural choice that adds complexity. This prompts the designer to explicitly rationalize trade-offs early in the implementation process.
@@ 7.2. Designing for Analytical Provenance to Support Resumption and Longer-term Usage
Reflexis is designed to bridge this gap by prioritizing a meaningful, auditable narrative of conceptual development over an exhaustive interaction log. We operationalize this via data-level provenance – which visualizes the structural evolution of analytic actions (e.g., splits, merges, renames) (Viégas et al., 2004), augmented by lightweight rationale provenance – the brief reflections and justifications linked to individual snippets. Mirroring narrative scaffolding work that suggests insight-linked provenance helps analysts act on reflections (Huang et al., 2026), our findings (RQ2) indicate that combining data-level provenance with lightweight rationales effectively balance utility and manageability. By transforming the codebook from a static list into a living history, the system shows potential as an external memory scaffold and is “a very useful way to think back.” This allows researchers to combat the “drift” and fatigue inherent in longitudinal work (O’Connor and Joffe, 2020). While longitudinal deployment is necessary to confirm long-term adoption, this immediate utility suggests that provenance views within Reflexis can facilitate re-immersion and transparent reporting, without the overhead of manual journaling.